# Single impurity atom embedded in a dipolar two-soliton molecule as a qubit

### v1

@@ Abstract
We consider a single impurity atom trapped in a double well (DW) potential created by a dipolar two-soliton molecule in a quasi-one-dimensional geometry. By solving the eigenvalue problem for the impurity atom in the DW potential, we find that its ground and first excited states are well separated from higher excited states. This allows it to be approximated by a desirable two-level quantum system. Numerical simulations of the Schrödinger equation, governing impurity atom, demonstrate periodic oscillations in the probability of finding the impurity confined either to the “left” or to the “right” side of the DW potential. An analytic expression for the coherent oscillations of the population imbalance between the two wells of the DW potential has been derived using the two-mode approximation. Theoretical predictions of the mathematical model are in good agreement with the results of numerical simulations. Potential usage of the developed setup as a physical realization of “qubit” has been discussed.
@@ I Introduction
The enhanced computational power of quantum computers will be primarily due to the use of quantum bits called qubit. Unlike its classical counterpart, called bit, which enables operations using only two digits “0” and “1”, qubits can exist in multiple states simultaneously, owing to the superposition principle of quantum mechanics, thus leading to an enormous increase in speed of operations. To date, the qubits have been realized on different physical platforms, such as Josephson junctions, quantum dots, nuclear magnetic resonances, trapped ions, neutral atoms in optical tweezers, cold atoms in optical lattice, coherent photons, etc. (see e.g. book nakahara-book ). Each of these realizations has distinct advantages and limitations, and lack of a universally accepted platform motivates the search for alternative realizations. Impurity atoms confined to matter-wave bright and dark solitons in the context of their qubit properties were investigated in Refs. shaukat2017 ; shaukat2019 . The key benefit of this approach is that the qubit is well protected from decoherence salerno2005 . Previous research primarily examined single hump matter-wave solitons with an embedded impurity atom. The energy levels of a particle in a single well potential, such as Gaussian nandi2010 or Pöschl-Teller type flugge-book ; senn1986 , are well known to be considerably spaced apart, making them less suitable for the functioning of a qubit. In contrast, double-well potentials exhibit closely spaced energy levels, making them suitable for qubit representation foot2011 . In this context, tunnel-coupled semiconductor double quantum dots were considered highly promising systems divincenzo2005 ; wiel2003 .
@@
Here, we propose a qubit realization using a single impurity atom embedded in a dipolar two-soliton molecule that forms a self-induced double well (DW) potential. This configuration enables a robust two-level qubit system, while the dipolar nonlocality offers in-situ tunability of the barrier height/width and energy splitting through the strength of inter-atomic interactions. It is important to note that matter wave solitons in Bose-Einstein condensates (BEC) with only contact atomic interactions do not support stable bound states resembling molecules. Therefore, ordinary BECs cannot provide a stable DW potential for the impurity atom.
@@
Our strategy in this work goes as follows: we first develop a variational approach to find the stationary state of the dipolar two-soliton molecule, then numerically calculate the energy spectrum of the impurity atom trapped in the molecular potential. In the framework of the two-mode approximation we derive the frequency of coherent oscillations of the population imbalance between the two wells of the DW potential. To validate our findings, the analytical predictions will be compared with numerical simulations. Finally, we discuss the potential usage of our results in the design a soliton-based qubit.
@@ II The wave profile of the two-soliton molecule
The model is described by a 1D nonlocal GPE supporting a two-soliton molecule and a linear Schrödinger equation for a single impurity atom embedded in the molecular potential
@@
The objective of this work is to develop an analytical framework for the “qubit problem”, based on the variational approach (VA). The validity of the VA for describing the properties of dipolar BEC over a wide range of experimental parameters was demonstrated in yi2001 ; olson2013 . It is evident from the basic model Eqs. (1)-(2) that the GPE is decoupled from the Schrödinger equation, while the impurity “feels” the dipolar BEC only as an external potential. Obviously, a single impurity atom cannot significantly distort the wave profile of the BEC, which consists of a large number of dipolar atoms. Similar models have been employed, and proven to be realistic, in previous publications devoted to BEC-based qubits shaukat2017 ; shaukat2019 ; salerno2005 .
@@
${\cal L}=\frac{i}{2}(\psi\psi^{\ast}_{t}-\psi^{\ast}\psi_{t})+\frac{1}{2}|\psi_{x}|^{2}-\frac{q}{2}|\psi|^{4}-\frac{g}{2}|\psi|^{2}\int\limits_{-\infty}^{\infty}R(x-x^{\prime})|\psi(x^{\prime},t)|^{2}dx^{\prime},$ (4)
@@
where $A(t)$, $a(t)$, $b(t)$ and $\varphi(t)$ are variational parameters, associated with the amplitude, width, chirp and phase, respectively. The norm $N=\int|\psi(x)|^{2}dx=A^{2}a^{3}\sqrt{\pi}/2$ is proportional to the number of atoms in the dipolar condensate and it represents the conserved quantity of GPE, (1). Substitution of the ansatz (5) and response function (3) into the Lagrangian density (4) and subsequent integration over the space variable $x$ yields the effective Lagrangian
@@
which is plotted in Fig. 1a for particular parameter settings. The minimum of the potential corresponds to the stationary width of the soliton molecule $a_{0}$, whose value can be found from the fixed point of Eq. (7), defined by $a_{tt}=0$, for particular parameter values $g$, NN and $w$. The separation between center-of-mass positions of two anti-phase bright solitons constituting the molecule is $\xi=4a_{0}/\sqrt{\pi}$. At larger separation ($a>a_{0}$) the solitons attract each other ($\partial U/\partial a>0$), and at smaller separation ($a<a_{0}$) they repel ($\partial U/\partial a<0$), therefore the effective potential $U(a)$ has a binding nature necessary for the formation of the stable molecule baizakov2019 . The wave profile of the dipolar two-soliton molecule, numerically constructed using the imaginary time propagation method chiofalo2000 applied to GPE (1), and as predicted by VA, is shown in Fig. 1b.
@@ IV Dynamics of the impurity atom in a double-well potential
$z(t)=\cos(\Omega t),\qquad\mbox{with}\qquad\Omega=\Delta/2.$ (20)
Figure 3 illustrates the real-time dynamics of the coupled system (1)–(2) for the parameter set $g=10$, $q=1$, $\gamma=3$, and $N=10$. Panel (a) shows the stationary double-well potential $|\psi(x,t)|^{2}$ created by the dipolar two-soliton molecule. As expected, the molecular density remains practically unchanged during the evolution, confirming that the impurity atom has a negligible back action on the condensate. Panel (b) presents the spatiotemporal evolution of the density $|\phi(x,t)|^{2}$, where the impurity periodically tunnels between the left and right wells of the self-induced potential, maintaining its localization and phase coherence over long time scales. Panel (c) displays the corresponding population imbalance
@@
Figure 3: (a) The DW potential $\sim|\psi(x,t)|^{2}$ created by the dipolar two-soliton molecule remains stationary as time advances. (b) The density plot for $|\phi(x,t)|^{2}$ shows regular oscillations of the probability of finding the impurity atom on the left/right well. (c) The corresponding population imbalance $z(t)=\int_{x<0}|\phi(x,t)|^{2}dx-\int_{x>0}|\phi(x,t)|^{2}dx$ demonstrates coherent Josephson-type oscillations. The numerically measured oscillation period $T_{\mathrm{meas}}=223.506$ agrees closely with the theoretical prediction $T_{\mathrm{pred}}=223.125$ obtained from the two-mode model [Eq. (20)]. The parameters used are $g=10$, $q=1$, $w=5$, $\gamma=3$, and $N=10$.
@@ V Tuning the qubit’s frequency
The ability to change the qubit’s frequency is crucial for its precise control and manipulation. Since the qubit’s frequency is determined by the energy difference $\Delta$ between its two lowest quantum states $\{\phi_{0},\phi_{1}\}$, the goal can be achieved by manipulating the underlying DW potential $V(x)\sim-|\psi(x)|^{2}$ in which the impurity atom is trapped. The DW potential can be modified by tuning the strength of atomic interactions in BEC, either short-range contact interactions $q$, or long-range dipolar interactions $g$. The variation of these parameters will result in corresponding variations of the distance $\sim a_{0}$ between the two wells of the DW potential, thus the desired energy splitting $\Delta=\mu_{1}-\mu_{0}$. When the two potential wells are sufficiently distant, the energy splitting disappears as the impurity atom’s tunneling probability approaches zero. In Fig. 4 we show the qubit’s frequency as a function of the strength of contact and dipolar atomic interactions. The curves demonstrate the tendency for the energy splitting $J=(\mu_{1}-\mu_{0})/2$ (proportional to the qubit’s oscillation frequency) to increase, as the two wells of the DW potential approach each other (width parameter $a_{0}$ decrease). Evidently, the effect of contact interactions is almost linear, while the long range dipolar interactions show hyperbolic behavior.
@@
Figure 4: The frequency of coherent oscillations of the impurity atom between the two wells of the double-well (DW) potential as a function of the interaction strengths in the dipolar condensate. (a) Dependence of the oscillation frequency on the contact interaction parameter $q$ for fixed $g=10$, $w=5$, $\gamma=3$, and $N=10$. (b) Dependence on the dipolar (nonlocal) interaction strength $g$ for fixed $q=1$, $w=5$, $\gamma=3$, and $N=5$. Blue circles represent the numerically obtained oscillation frequencies, while the yellow dashed curves correspond to the analytical predictions based on the variationally obtained width parameter $a_{0}$ from Eq. (8).
@@
Figure 5: Left panel: The Bloch sphere represents a qubit state. The north pole corresponds to the ground state $\phi_{0}$ and the south pole corresponds to the first excited state $\phi_{1}$. An arbitrary superposition of these basis states can be obtained by parametrization $\phi=\cos(\theta/2)|\phi_{0}\rangle+e^{i\varphi}\sin(\theta/2)|\phi_{1}\rangle$. Right panel: The Bloch components $r_{x}\simeq\cos(\Delta t)$, $r_{y}\simeq\sin(\Delta t)$, $r_{z}\simeq 0$ show circular equatorial motion (upper). The purity of the two-level system calculated as ${\rm Tr}(\rho^{2})$, with $\rho$ being the density matrix, remains very close to unity indicating negligible de-coherence (lower). The parameters used are $g=10$, $q=1$, $w=5$, $\gamma=3$, and $N=10$.
VI Conclusions

### v2

@@ Abstract
We consider a single impurity atom trapped in a double well (DW) potential created by a dipolar two-soliton molecule in a quasi-one-dimensional geometry. By solving the eigenvalue problem for the impurity atom in the DW potential, we find that its ground and first excited states are well separated from higher excited states. This allows it to be approximated by a desirable two-level quantum system. Numerical simulations of the Schrödinger equation, governing impurity atom, demonstrate periodic oscillations in the probability of finding the impurity confined either to the “left” or to the “right” well of the DW potential. An analytic expression for the coherent oscillations of the population imbalance between the two wells of the DW potential has been derived using the two-mode approximation. Theoretical predictions of the mathematical model are in good agreement with the results of numerical simulations. Potential usage of the developed setup as a physical realization of “qubit” has been discussed.
@@ I Introduction
The enhanced computational power of quantum computers will be primarily due to the use of quantum bits called qubit. Unlike its classical counterpart, called bit, which enables operations using only two digits “0” and “1”, qubits can exist in multiple states simultaneously, owing to the superposition principle of quantum mechanics, thus leading to an enormous increase in speed of operations. To date, the qubits have been realized on different physical platforms, such as Josephson junctions, quantum dots, nuclear magnetic resonances, trapped ions, neutral atoms in optical tweezers, cold atoms in optical lattice, coherent photons, etc. (see e.g. book nakahara-book ). Each of these realizations has distinct advantages and limitations, and lack of a universally accepted platform motivates the search for alternative realizations. Impurity atoms confined to matter-wave bright and dark solitons in the context of their qubit properties were investigated in Refs. shaukat2017 ; shaukat2019 . The key benefit of this approach is that the qubit is well protected from decoherence salerno2005 . Previous research primarily examined single hump matter-wave solitons with an embedded impurity atom. The energy levels of a particle in a single well potential, such as Gaussian nandi2010 or Pöschl-Teller type flugge-book ; senn1986 , are well known to be considerably spaced apart, making them less suitable for the functioning of a qubit with a desired frequency and tunability. In contrast, double-well potentials exhibit closely spaced energy levels, making them suitable for qubit representation foot2011 . In this context, tunnel-coupled semiconductor double quantum dots were considered highly promising systems divincenzo2005 ; wiel2003 .
@@
Here, we propose a qubit realization using a single impurity atom embedded in a dipolar two-soliton molecule that forms a self-induced double well (DW) potential. This configuration enables a robust two-state quantum system, while the dipolar nonlocality offers in-situ tunability of the barrier height/width and energy splitting (qubit frequency) through the strength of inter-atomic interactions. The present model illustrates a qubit that is delocalized in space, with its computational basis states determined by the presence or absence of an impurity atom in one of the two potential wells of a double-well trap. A notable advantage of this approach is the ability to create highly entangled states involving multiple qubits mompart2003 . It is important to note that matter wave solitons in Bose-Einstein condensates (BEC) with only contact atomic interactions do not support stable bound states resembling molecules. Therefore, ordinary BECs cannot provide a stable DW potential for the impurity atom. In a different context, the applications of BECs for quantum computation and information processing have been reported in the literature byrns2012 ; xu2022 ; boudjemaa2025 ; barshilia2024 ; shaukat2020 ; shaukat2020a .
@@
Our strategy in this work goes as follows: we first develop a variational approach to find the stationary state of the dipolar two-soliton molecule, then numerically calculate the energy spectrum of the impurity atom trapped in the molecular potential. In the framework of the two-mode approximation we derive the frequency of coherent oscillations of the population imbalance between the two wells of the DW potential. To validate our findings, the analytical predictions will be compared with numerical simulations. Finally, we discuss the potential usage of our results in the design of a soliton-based qubit.
@@ II The wave profile of the two-soliton molecule
The essential features of the present system can be understood by considering the model in a quasi-one-dimensional geometry. Reducing the coupled 3D Gross-Pitaevskii equations (GPE) to a quasi-1D form was elaborated in numerous publications, among which we refer to a recent paper devoted to binary dipolar BECs adhikari2024 . The results of this work suggest that the influence of the minority component (single non-dipolar atom) on the majority component (soliton molecule consisting of ${\cal N}\sim 10^{3}\div 10^{4}$ dipolar atoms) is proportional to $1/{\cal N}$. Consequently, the effect of the impurity on the soliton molecule can be considered negligible. In these conditions, the model is described by a 1D nonlocal GPE supporting a two-soliton molecule and a linear Schrödinger equation for a single impurity atom embedded in the molecular potential
@@
The objective of this work is to develop an analytical framework for the “qubit problem”, based on the variational approach (VA). The validity of the VA for describing the properties of dipolar BEC over a wide range of experimental parameters was demonstrated in yi2001 ; olson2013 . It is evident from the basic model Eqs. (1)-(2) that the GPE is decoupled from the Schrödinger equation, while the impurity “feels” the dipolar BEC only as an external potential. Obviously, a single impurity atom cannot significantly distort the wave profile of the BEC, which consists of a large number of dipolar atoms wenzel2018 . Similar models have been employed, and proven to be realistic, in previous publications devoted to BEC-based qubits shaukat2017 ; shaukat2019 ; salerno2005 .
@@
${\cal L}=\frac{i}{2}(\psi\psi^{\ast}_{t}-\psi^{\ast}\psi_{t})+\frac{1}{2}|\psi_{x}|^{2}-\frac{q}{2}|\psi|^{4}-\frac{g}{2}|\psi|^{2}\int\limits_{-\infty}^{\infty}R(x-x^{\prime})|\psi(x^{\prime},t)|^{2}dx^{\prime}.$ (4)
@@
where $A(t)$, $a(t)$, $b(t)$ and $\varphi(t)$ are variational parameters, associated with the amplitude, width, chirp and phase, respectively. The norm $N=\int|\psi(x)|^{2}dx=A^{2}a^{3}\sqrt{\pi}/2$ is proportional to the number of atoms in the dipolar condensate and it represents the conserved quantity of GPE (1). The main advantage of the trial function (5) over alternative two anti-phase Gussians otajonov2017 is that it leads to simpler variational equations, although its validity is limited to small deviations of solitons from their equilibrium positions.
Substitution of the ansatz (5) and response function (3) into the Lagrangian density (4) and subsequent integration over the space variable $x$ yields the effective Lagrangian
@@
which is plotted in Fig. 1a for particular parameter settings. The minimum of the potential corresponds to the stationary width of the soliton molecule $a_{0}$, whose value can be found from the fixed point of Eq. (7), defined by $a_{tt}=0$, for particular parameter values $q$, $g$, NN and $w$. The separation between center-of-mass positions of two anti-phase bright solitons constituting the molecule is $\xi=4a_{0}/\sqrt{\pi}$. At larger separation ($a>a_{0}$) the solitons attract each other ($\partial U/\partial a>0$), and at smaller separation ($a<a_{0}$) they repel ($\partial U/\partial a<0$), therefore the effective potential $U(a)$ has a binding nature necessary for the formation of the stable molecule baizakov2019 . The wave profile of the dipolar two-soliton molecule, numerically constructed using the imaginary time propagation method chiofalo2000 applied to GPE (1), and as predicted by VA, is shown in Fig. 1b.
@@
The stability of the soliton molecule has been verified by real-time GPE evolution of the spatially perturbed initial wave profile. During time propagation, the soliton molecule sheds the perturbation as linear waves, which are absorbed at the integration domain boundaries and quickly acquires a smooth profile, thereby demonstrating its stability.
@@ IV Dynamics of the impurity atom in a double-well potential
$z(t)=\cos(2\Omega t),\qquad\mbox{with}\qquad\Omega=\Delta/2.$ (20)
Figure 3 illustrates the real-time dynamics of the coupled system (1)–(2) for the parameter set $q=1$, $g=10$, $\gamma=3$, and $N=10$. Panel (a) shows the stationary double-well potential $|\psi(x,t)|^{2}$ created by the dipolar two-soliton molecule. As expected, the molecular density remains practically unchanged during the evolution, confirming that the impurity atom has a negligible back action on the condensate. Panel (b) presents the spatiotemporal evolution of the density $|\phi(x,t)|^{2}$, where the impurity periodically tunnels between the left and right wells of the self-induced potential, maintaining its localization and phase coherence over long time scales. Panel (c) displays the corresponding population imbalance
@@
Figure 3: (a) The DW potential $\sim|\psi(x,t)|^{2}$ created by the dipolar two-soliton molecule remains stationary as time advances. (b) The density plot for $|\phi(x,t)|^{2}$ shows regular oscillations of the probability of finding the impurity atom on the left/right well. (c) The corresponding population imbalance $z(t)=\int_{x<0}|\phi(x,t)|^{2}dx-\int_{x>0}|\phi(x,t)|^{2}dx$ demonstrates coherent Josephson-type oscillations. The numerically measured oscillation period $T_{\mathrm{meas}}=223.506$ agrees closely with the theoretical prediction $T_{\mathrm{pred}}=223.125$ obtained from the two-mode model [Eq. (20)]. The parameters used are $q=1$, $g=10$, $w=5$, $\gamma=3$, and $N=10$.
@@ V Tuning the qubit’s frequency
The ability to change the qubit’s frequency is crucial for its precise control and manipulation. Since the qubit’s frequency is determined by the energy difference $\Delta$ between its two lowest quantum states $\{\phi_{0},\phi_{1}\}$, the goal can be achieved by manipulating the underlying DW potential $V(x)\sim-|\psi(x)|^{2}$ in which the impurity atom is trapped. The DW potential can be modified by tuning the strength of atomic interactions in BEC, either short-range contact interactions $q$, or long-range dipolar interactions $g$. The variation of these parameters will result in corresponding variations of the distance $\sim a_{0}$ between the two wells of the DW potential, thus the desired energy splitting $\Delta=\mu_{1}-\mu_{0}$. When the two potential wells are sufficiently distant, the energy splitting disappears as the impurity atom’s tunneling probability approaches zero. In Fig. 4 we show the qubit’s frequency as a function of the strength of contact and dipolar atomic interactions. The curves demonstrate the tendency for the energy splitting $J=(\mu_{1}-\mu_{0})/2$ (proportional to the qubit’s oscillation frequency) to increase, as the two wells of the DW potential approach each other (width parameter $a_{0}$ decrease). Evidently, the effect of contact interactions is almost linear, while the long range dipolar interactions show hyperbolic behavior. The system’s tunability facilitates controlled interactions between neighboring qubits in an array, enabling the generation of entangled states that are essential for robust quantum computations and effective error correction schemes.
@@
Figure 4: The frequency of coherent oscillations of the impurity atom between the two wells of the double-well potential as a function of the interaction strengths in the dipolar condensate. (a) Dependence of the oscillation frequency on the contact interaction parameter $q$ for fixed $g=10$, $w=5$, $\gamma=3$, and $N=10$. (b) Dependence on the dipolar (nonlocal) interaction strength $g$ for fixed $q=1$, $w=5$, $\gamma=3$, and $N=5$. Blue circles represent the numerically obtained oscillation frequencies, while the yellow dashed curves correspond to the analytical predictions based on the variationally obtained width parameter $a_{0}$ from Eq. (8).
@@
Figure 5: Left panel: The Bloch sphere illustrates the qubit state generated from direct numerical simulations as described in the main text. The north pole corresponds to the ground state $\phi_{0}$ and the south pole corresponds to the first excited state $\phi_{1}$. An arbitrary superposition of these basis states can be obtained by parametrization $\phi=\cos(\theta/2)|\phi_{0}\rangle+e^{i\varphi}\sin(\theta/2)|\phi_{1}\rangle$. Right panel: The Bloch components $r_{x}\simeq\cos(\Delta t)$, $r_{y}\simeq\sin(\Delta t)$, $r_{z}\simeq 0$ show circular equatorial motion (upper). The purity of the two-level system calculated as ${\rm Tr}(\rho^{2})$, with $\rho$ being the density matrix, remains very close to unity indicating negligible de-coherence (lower). The parameters used are $q=1$, $g=10$, $w=5$, $\gamma=3$, and $N=10$.
VI Prospects of experimental realization
The concept of this work is based on the properties of a two-state quantum system formed by an impurity atom within a double-well potential, induced by a dipolar two-soliton molecule. Although the feasibility of creating soliton molecules in dipolar BECs has been theoretically demonstrated in several publications lakomy2012 ; edmonds2017 , their experimental realization has not yet been reported. Meanwhile, soliton molecules, consisting of stable bound states of two or more bright solitons, have been experimentally observed in dispersion-managed optical fibers stratmann2005 ; rohrmann2013 and mode-locked fiber lasers tang2001 ; liu2022 . Advances in quantum technologies, based on a soliton-impurity system, may inspire experiments to create dipolar soliton molecules.
[…]
Immersing an impurity atom into a Bose-Einstein condensate is an advanced experimental procedure typically performed using techniques like laser cooling and magnetic/optical trapping, followed by precise control over the impurity’s position and interaction strength mayer2019 . We analyze the tunneling of an impurity atom through a barrier in a double-well potential using the two-state approximation given by Eq. (12). If the impurity is initially placed in a left well, represented as $\phi(x,0)=\phi_{L}(x)$, the probability density $|\phi(x,t)|^{2}$ will oscillate periodically over time, with a period given by $T=2\pi\hbar/\Delta$. Here, $\Delta=\mu_{1}-\mu_{0}$ denotes the energy difference between the first excited state and the ground state. This difference is typically interpreted as the energy splitting of the ground-state level in a single well, which occurs due to the coupling between the two wells. In our calculations, the length scale for the impurity atom is considered to be $\sim a_{s}$, which yields the characteristic frequency $\omega_{0}=\hbar/m_{Rb}a_{s}^{2}\simeq 26.5$ MHz for $m_{Rb}=1.4\times 10^{-25}$ kg. Therefore the qubit frequency shown in Fig. 3c is $\omega=(2\pi/T_{meas})\cdot\omega_{0}\simeq 750$ kHz.
In this study, we considered the functioning of the proposed qubit at the proof-of-principle level. Further research will focus on exploring the influence of adverse effects, such as thermal noise, phonon scattering, fluctuations in the soliton molecule, etc.
VII Conclusions